Light falls on a magnesium surface. The work function of magnesium is 3.68 eV.
Find the minimum frequency of the light source required to eject the
photoelectrons. Can the photoelectrons be ejected by increasing the intensity of
the light source? Why?
1
Expert's answer
2019-10-23T09:53:21-0400

The threshold frequency for the magnesium


"f_0=\frac{W}{h}=\frac{3.68\:\rm eV}{4.135\times 10^{-15}\:\rm eV\cdot s}=8.9\times 10^{14}\: \rm Hz"

The increasing of the the intensity of the light source causes increasing of number of ejected photoelectrons.